# Sample-Efficiency in Multi-Batch Reinforcement Learning: The Need for Dimension-Dependent Adaptivity

### v1

@@ E.1 MDP Class
The BPI and PE problems used in the proof of Theorem 4.5 are characterised by a class of MDPs $\mathcal{M}$ and target policies $\Pi$. In this section, we define the class of MDPs $\mathcal{M}$. All MDPs $M\in\mathcal{M}$ in the class share the same state-space $\mathcal{S}$, action space $\mathcal{A}$, feature map and target policy $\pi$ (for the PE problem) but differ in the transition function $p_{M}$ and reward function $r_{M}$. This class of MDPs is similar to the one used by Zanette (2021) in the proof of their Theorem 3. Our construction differs in the transition functions which are defined in Appendix E.2.
@@ E.2 Instance of the class
s.t $\dim B_{K}>0$ and fix some $w\in B_{K}\cap\partial\mathcal{B}$. Set $B_{K+1}=\langle w\rangle$. Let $H_{w}=\{B_{1},...,B_{K},B_{K+1}\}$ denote the set of nested subspaces (including $\langle w\rangle$). This is not defined as an ordered set, but for notational purposes the order can always be recovered since the sequence must be nested.
@@
See Appendix B for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$. We sometimes use a ${{w,\pm}}$ subscript to refer to both MDPs simultaneously.
Transition Function: The transition function for an MDP $M\in\mathcal{M}$ depends on the sequence of nested subspaces $H_{w}$ but not on the sign $\pm$. The transition function is therefore the same for $M_{H_{w},+}$ and $M_{H_{w},-}$. If $A$ is a subspace of $\mathbb{R}^{d}$, let $\text{proj}_{A}(x)$ denote the orthogonal projection of $x$ onto $A$. See Appendix B for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$ and sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. The successor state of a state-action pair $(s,a)$ is deterministic and only depends on the chosen action (and not the current state), so we will denote the unique successor state when taking action $a$ by $s^{+}_{H_{w}}(a)=a$, which is defined as:
@@
The transition function is defined for all $a\in\mathcal{A}=\bigcup_{s\in\mathcal{S}}\mathcal{A}_{s}=\mathcal{B}$:
$=B_{0}$
$=\mathcal{B}.$
and the pair-wise intersections are all empty (so they form a partition). The successor states are also well-defined:
  • For $k=0,...,K$, if $a\in\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k})\setminus\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$, then $a\notin\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$ and from Definition B.2, this means that for any $u\in B_{k+1}$, $|u^{T}s|/\|u\|_{2}\|s\|_{2}\leq\gamma$. Denoting $x=\text{proj}_{B_{k+1}}(a)\in B_{k+1}$,
    $\|x\|_{2}=\frac{x^{T}x}{\|x\|_{2}}=\frac{|x^{T}a|}{\|x\|_{2}}\leq\gamma\|a\|_{2}\leq\gamma,$
    meaning that $\frac{1}{\gamma}\text{proj}_{B_{k+1}}(a)\in\mathcal{B}$.
  • If $a\in\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{K+1})\setminus(\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w))$, , then $a\notin(\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w))$ and from Definition B.1, this means that $|w^{T}s|\leq\gamma$. Denoting $x=\text{proj}_{B_{K+1}}(a)\in B_{K+1}$,
    $\|x\|_{2}=\frac{x^{T}x}{\|x\|_{2}}=\frac{|x^{T}a|}{\|x\|_{2}}\leq\gamma,$
    meaning that $\frac{1}{\gamma}\text{proj}_{B_{K+1}}(a)\in\mathcal{B}$.
Note that the starting state $\bar{s}$ is not a successor state so a policy trajectory can never return to $\bar{s}$.
@@
E.3 Realizability
We show that the action-value of any policy (not just the target policy) for any MDP $M\in\mathcal{M}$ can be linearly represented with the feature map defined in Appendix E.1. The action-value of a policy $\pi$ only depends on $w$ and the sign $\pm$ of the reward function of the MDP, the sequence of nested subspace does not matter beyond $w$.
Lemma E.1 (Realizability).
For any $w\in\partial\mathcal{B}$ and sequence of nested subspaces $H_{w}$ satisfying the construction from Appendix E.2, let $Q^{\pi}_{H_{w},+}$ and $Q^{\pi}_{H_{w},-}$ be the action-value functions of an arbitrary policy $\pi$ on $M_{H_{w},+}$ and $M_{H_{w},-}$, respectively. Then it holds that
@@ E.3 Realizability
For all cases, $Q$ satisfies the Bellman evaluation equations, so it is the fixed point of the Bellman evaluation operator. In particular, it is the action-value of the policy $\pi$ on $M_{H_{w},+}$.
Now consider $M_{w,-}$, for which the argument is identical. Set $Q(s,a)=-\phi(s,a)^{T}w$. We will show that $Q$ satisfies the Bellman evaluation equations for $\pi$ at all state-actions pairs, which will imply that $Q(s,a)=Q^{\pi}_{H_{w},-}(s,a)$. Apply $\mathcal{T}^{\pi}_{M_{H_{w},-}}$ to $Q$ at $(s,a)$:
$\mathcal{T}^{\pi}_{M_{H_{w},-}}(Q)(s,a)$ $=r_{H_{w},-}(s,a)+\gamma Q(s^{+}_{H_{w}}(a),\pi(s^{+}_{H_{w}}(a))).$
We consider the RHS of the above for all possible cases.
$r_{H_{w},-}(s,a)+\gamma Q(s^{+}_{H_{w}}(a),\pi(s^{+}_{H_{w}}(a)))$ $=0+\gamma Q(s^{+}_{H_{w}}(a),s^{+}_{H_{w}}(a))$
$=-\gamma\phi(s^{+}_{H_{w}}(a),s^{+}_{H_{w}}(a))^{T}w$
$=-\gamma s^{+}_{H_{w}}(a)^{T}w$
$=-\gamma\frac{1}{\gamma}\text{proj}^{T}_{B_{k+1}}(a)w$
$=-\text{proj}^{T}_{B_{k+1}}(a)w.$
However, recall that $w\in B_{k+1}$. Since $\text{proj}_{B_{k+1}}(a)$ is the orthogonal projection of $a$: $w^{T}\text{proj}_{B_{k+1}}(a)=w^{T}a$. Plugging into the above, we have: $\mathcal{T}^{\pi}_{M_{H_{w},+}}(Q)(s,a)=-a^{T}w=-\phi(s,a)^{T}w=-Q(s,a)$, which satisfies the Bellman evaluation equation.
$r_{H_{w},-}(s,a)+\gamma Q(s^{+}_{H_{w}}(a),\pi(s^{+}_{H_{w}}(a)))$ $=0+\gamma Q(s^{+}_{H_{w}}(a),s^{+}_{H_{w}}(a))$
$=-\gamma\phi(s^{+}_{H_{w}}(a),s^{+}_{H_{w}}(a))^{T}w$
$=-\gamma s^{+}_{H_{w}}(a)^{T}w$
$=-\gamma\frac{1}{\gamma}\text{proj}^{T}_{B_{K+1}}(a)w$
$=-\text{proj}^{T}_{B_{K+1}}(a)w$
$=-a^{T}w=-\phi(s,a)^{T}w=Q(s,a),$
which satisfies the Bellman evaluation equation.
$r_{H_{w},-}(s,a)+\gamma Q(s^{+}_{H_{w}}(a),\pi(s^{+}_{H_{w}}(a)))$ $=-(1-\gamma)a^{T}w+\gamma Q(a,a)$
$=-(1-\gamma)a^{T}w-\gamma\phi(a,a)^{T}w$
$=-(1-\gamma)a^{T}w-\gamma a^{T}w$
$=-a^{T}w=-\phi(s,a)^{T}w=Q(s,a),$
which satisfies the Bellman evaluation equation.
For all cases, $Q$ satisfies the Bellman evaluation equations, so it is the fixed point of the Bellman evaluation operator. In particular, it is the action-value of the policy $\pi$ on $M_{H_{w},-}$. ∎
E.4 Proof of Theorem 4.5
Consider the MDP class described in Appendix E.1 and Appendix E.2. First, we know from Lemma E.1 that all instances of the BPI and PE problem characterised by the MDP class $\mathcal{M}$ and target policies $\Pi$ satisfy Assumption 4.2 ($Q^{\pi}$ is realizable for every $\pi$) with the feature map $\phi(\cdot,\cdot)$ defined in Appendix E.1. The proof follows the same reasoning as in Appendix D.4. We consider policy-free queries.
@@
The learner is sample-efficient so $n$ is polynomial in $d$. Specifically, there exists some constant $\alpha>0$ and some integer $T$ such that $n\leq\alpha d^{T}$. Consider $N\in\mathbb{N}$ s.t $2^{N}\leq d<2^{N+1}$ and set $d_{+}=2^{N}$. Note that $d_{+}\geq d/2$. Fix $W=exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{K}})$. If $n\leq W$, we can show the learner cannot solve the PE problem. Consider both cases:
E.4.1 Case 1
Suppose $W<n$, then
$W<n$ $\implies W<\alpha d^{T}$
$\implies\frac{g(\gamma)}{8}d_{+}^{1/4^{K}}<\log(\alpha d^{T})$
$\implies(d/2)^{1/4^{K}}<\frac{8}{g(\gamma)}\log(\alpha d^{T})$
E.4.2 Case 2
$n\leq W\implies n_{k}\leq\exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{k}})$ for $k=1,...,K$. We inductively define the sequence of nested subspaces $B_{K}\subset...\subset B_{1}$ characterising $w$ ($\in B_{K}$) and used in the successor state function $s^{+}_{M}$ s.t for $k=1,...,K$, $\dim B_{k}=2^{\lceil N/4^{k}\rceil}\geq 2^{8}$ and
@@ E.4.2 Case 2
Proof of existence of nested subspaces:
We proceed by induction. Let $B_{K}\subset...\subset B_{1}$ be an arbitrary sequence of subspaces and $w\in B_{K}\cap\mathcal{B}$ arbitrary. We will define these but for now consider the successor state function $s^{+}_{H_{w}}$ (see Appendix E.2) with $H_{w}=\{\langle w\rangle,B_{K},...,B_{1}\}$.
so by Lemma C.3 there exists a subspace $H\in\mathcal{G}_{2^{\lceil N/4\rceil},d_{+}}(\mathbb{R})$ of dimension $2^{\lceil N/4\rceil}$ ($\geq d_{+}^{1/4}$) such that
$\{(r_{w,\pm}(s,a),s^{+}_{H_{w}}(a)):a\in A_{1}\}.$
Not having fixed $B_{2},...,B_{K},w$ does not cause problems with the feedback even though $s^{+}_{H_{w}}$ and $r_{w,\pm}$ depend on them since the feedback of $A_{1}$ only depends on $B_{1}$: the learner only observes the projection of the actions in $A_{1}$ into $B_{1}$,
$\implies$ $\forall a\in\widebar{A}_{1},\qquad s^{+}_{H_{w}}(a)=\frac{1}{\gamma}\text{proj}_{B_{1}}(a),\qquad r_{w,\pm}(s,a)=0.$
Inductive Step: Suppose that at round $k$, there exists a sequence of nested subspaces $B_{k}\subset...\subset B_{1}$ used by the successor state function $s^{+}_{H_{w}}$ s.t for $i=1,...,k$, $\dim B_{i}=2^{\lceil N/4^{i}\rceil}\geq 2^{8}$ and
$\implies$ $\forall a\in{A}_{k},\qquad s^{+}_{H_{w}}(a)=\frac{1}{\gamma}\text{proj}_{B_{j}}(a)\qquad\text{ for some }j\leq k,\qquad r_{w,\pm}(s,a)=0.$
Then we know that the number of those queries that are in $B_{k}$ is also less than $\Big{(}\frac{d_{k}}{2}\Big{)}^{\frac{1}{8}g(\gamma)d_{k}^{1/4}}$. Therefore by Lemma C.3, there exists a subspace $H\subset B_{k}$ of dimension $2^{\lceil\lceil N/4^{k}\rceil/4\rceil}\geq d_{k+1}=2^{\lceil N/4^{k+1}\rceil}$ (can reduce the dimension if they do not match, removing dimensions will not add queries to $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$) such that
$\{(r_{w,\pm}(s,a),s^{+}_{H_{w}}(a)):a\in A_{k+1}\}.$
Not having fixed $B_{k+2},...,B_{K},w$ does not cause problems with the feedback even though $s^{+}_{H_{w}}$ and $r_{w,\pm}$ depend on them since the feedback observed up to this round of $\widebar{A}_{k+1}$ only depends on $B_{1},...,B_{k+1}$: the learner only observes the projection of the actions in $\widebar{A}_{k+1}$ into $B_{1},...,B_{k+1}$,
$\forall a\in{A}_{k+1},\qquad a\notin\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$
$\implies$ $\forall a\in{A}_{k},\qquad s^{+}_{H_{w}}(a)=\frac{1}{\gamma}\text{proj}_{B_{j}}(a)\qquad\text{ for some }j\leq k+1,\qquad r_{w,\pm}(s,a)=0.$
We remark Lemma C.3 establishes the existence of a subspace in $\mathcal{G}_{2^{\lceil\lceil N/4^{k}\rceil/4\rceil},d_{k}}(\mathbb{R})$ where the ambient space is $\mathbb{R}^{d_{k}}$ instead of $\mathbb{R}^{d}$. Any $d_{k}$-dimensional subspace of $\mathbb{R}^{d}$ is isomorphic to $\mathbb{R}^{d_{k}}$, so we can consider $B_{k}$, project the points of $A_{k+1}$ into $B_{k}$, apply Lemma C.3 to get $H$ within $B_{k}$, and extend $H$ to be defined in $\mathbb{R}^{d}$. We omit these steps here for clarity but detail them fully in Appendix D.4.3. This also holds for the base case (where $d_{+}$ is used as the ambient dimension instead of $d$)
When applying Lemma C.3 with ambient dimension $d_{k}=2^{\lceil N/4^{k}\rceil}$, we require $\lceil N/4^{k}\rceil\geq 8$. Since $d_{k}\geq d_{+}^{1/4^{k}}$ and $d_{+}\geq d/2$, it is enough to have
$\impliedby K\leq c\log\log d,$
$K>c\log\log d.$ (4)
End of proof of existence of nested subspaces.
@@
If the learner outputs a policy taking action $a$ s.t $a^{T}w>0$, it will incur an error greater than $1$ on $M_{H_{w},-}$ and similarly for $M_{H_{w},+}$ if it produces an action $a$ s.t $a^{T}w\leq 0$. Even if it randomizes between both, with probability at least $1/2$ it will incur an error of at least $1$ on one of the MDPs.
@@
Appendix F Proofs for Fully-Adaptive Setting
F.1 Proof of Theorem A.2
@@ Appendix F Proofs for Fully-Adaptive Setting
Step 1: The learner selects an arbitrary query $(s_{1},a_{1})$ s.t $\|\phi(s_{1},a_{1})\|_{2}=1$ (possible by Assumption A.1). The learner receives from the environment the reward $r_{M}(s_{1},a_{1})$, the transition function $p_{M}(\cdot|s_{1},a_{1})$ and evaluations of the target policy $\pi_{M}$ for all states in the support of the transition function $p_{M}(\cdot|s_{1},a_{1})$.
@@
The feedback to the learner up to round $k$ means the learner has access to $v_{1},...,v_{k-1}$. This together with Assumption A.1 ensures the query-choice is possible.
@@
$\exists(\alpha_{i})_{i=1}^{k}\quad\text{ s.t one of them is non-zero and }\qquad\sum\limits_{i=1}^{k}\alpha_{i}v_{i}=0.$
Let $j$ be the largest index s.t $\alpha_{j}\neq 0$, then
@@
F.2 Proof of Theorem A.3
F.2.1 MDP Class
We consider the same MDP class as for Theorem E - see Appendix E.1.
F.2.2 Instance of the class
@@
s.t $\dim B_{k}=d-k$. Since $\dim B_{d-1}=1$, there is some $w\in B_{K}\cap\partial\mathcal{B}$ s.t. $B_{d-1}=\langle w\rangle$. Let $H_{w}=\{B_{1},...,B_{d-1}\}$ denote the set of nested subspaces. Every MDP $M\in\mathcal{M}$ is fully characterised by the sequence of subspaces $H_{w}$ and a sign $+$ or --. Hence, they are denoted by $M_{H_{w},+}$ or $M_{H_{w},-}$.
@@
See Appendix B for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$.
Transition Function: The transition function for an MDP $M\in\mathcal{M}$ depends on the sequence of nested subspaces $H_{w}$ but not on the sign $\pm$. The transition function is therefore the same for $M_{H_{w},+}$ and $M_{H_{w},-}$. If $A$ is a subspace of $\mathbb{R}^{d}$, let $\text{proj}_{A}(x)$ denote the orthogonal projection of $x$ onto $A$. See Appendix B for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$ and sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. The successor state of a state-action pair $(s,a)$ is deterministic and only depends on the chosen action (and not the current state), so we will denote the unique successor state when taking action $a$ by $s^{+}_{H_{w}}(a)=a$, which is defined as:
@@
This is defined in the same way as in the proof of Theorem E - we refer the reader to Appendix E.1 for an explanation of why this definition is well defined.
@@
F.2.3 Proof of Theorem A.3
Consider the MDP class described in Appendix F.2.1 and Appendix F.2.2. First, we know from Lemma E.1 that all instances of the BPI and PE problem characterised by the MDP class $\mathcal{M}$ and target policies $\Pi$ satisfy Assumption 4.2 ($Q^{\pi}$ is realizable for every $\pi$) with the feature map $\phi(\cdot,\cdot)$ defined in Appendix E.1.
@@
In Appendix F.2.2, we fixed $\dim B_{k}=d-k$. If $a_{1},...,a_{k}$ are not linearly independent, the dimension of $B_{k}$ may be greater than $d-k$. In this case, we just restrict $B_{k}$ arbitrarily such that the sequence remains nested. In particular, we have $\dim B_{d-1}=1$ and $B_{d-1}=\langle w\rangle$ for some $w\in\partial\mathcal{B}$. Let $H_{w}=\{B_{1},...,B_{d-1}\}$. Consider the BPI and PE problem instances with MDPs $M_{H_{w},+}$ and $M_{H_{w},-}$. The transition function is the same on $M_{H_{w},+}$ and $M_{H_{w},-}$.
@@
If the learner outputs a policy taking action $a$ s.t $a^{T}w>0$, it will incur an error greater than $1$ on $M_{H_{w},-}$ and similarly for $M_{H_{w},+}$ if it produces an action $a$ s.t $a^{T}w\leq 0$. Even if it randomizes between both, with probability at least $1/2$ it will incur an error of at least $1$ on one of the MDPs.

### v2

@@ F.1 MDP Class
The BPI and PE problems used in the proof of Theorem 4.5 are characterised by a class of MDPs $\mathcal{M}$ and target policies $\Pi$. In this section, we define the class of MDPs $\mathcal{M}$. All MDPs $M\in\mathcal{M}$ in the class share the same state-space $\mathcal{S}$, action space $\mathcal{A}$, feature map and target policy $\pi$ (for the PE problem) but differ in the transition function $p_{M}$ and reward function $r_{M}$. This class of MDPs is similar to the one used by Zanette (2021) in the proof of their Theorem 3. Our construction differs in the transition functions which are defined in Appendix F.2.
@@ F.2 Instance of the class
s.t. $\dim B_{K}>0$ and fix some $w\in B_{K}\cap\partial\mathcal{B}$. Set $B_{K+1}=\langle w\rangle$. Let $H_{w}=\{B_{1},...,B_{K},B_{K+1}\}$ denote the set of nested subspaces (including $\langle w\rangle$). This is not defined as an ordered set, but for notational purposes the order can always be recovered since the sequence must be nested.
@@
See Appendix C for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$. We sometimes use a ${{w,\pm}}$ subscript to refer to both MDPs simultaneously.
Transition Function: The transition function for an MDP $M\in\mathcal{M}$ depends on the sequence of nested subspaces $H_{w}$ but not on the sign $\pm$. The transition function is therefore the same for $M_{H_{w},+}$ and $M_{H_{w},-}$. If $A$ is a subspace of $\mathbb{R}^{d}$, let $\text{proj}_{A}(x)$ denote the orthogonal projection of $x$ onto $A$. See Appendix C for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$ and sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. The successor state of a state-action pair $(s,a)$ is deterministic and only depends on the chosen action (and not the current state), so we will denote the unique successor state when taking action $a$ by $s^{+}_{H_{w}}(a)=a$, which is defined as:
@@
Note that the starting state $\bar{s}$ is not a successor state so a policy trajectory can never return to $\bar{s}$. Showing that the transition function and successor states are well defined follows the same steps as showing the target policy is well defined in the proof of Theorem 4.4 in Section E.2.
@@
F.3 Realizability
We show that the action-value of any policy (not just the target policy) for any MDP $M\in\mathcal{M}$ can be linearly represented with the feature map defined in Appendix F.1. The action-value of a policy $\pi$ only depends on $w$ and the sign $\pm$ of the reward function of the MDP, the sequence of nested subspace does not matter beyond $w$.
Lemma F.1 (Realizability).
For any $w\in\partial\mathcal{B}$ and sequence of nested subspaces $H_{w}$ satisfying the construction from Appendix F.2, let $Q^{\pi}_{H_{w},+}$ and $Q^{\pi}_{H_{w},-}$ be the action-value functions of an arbitrary policy $\pi$ on $M_{H_{w},+}$ and $M_{H_{w},-}$, respectively. Then it holds that
@@ F.3 Realizability
For all cases, $Q$ satisfies the Bellman evaluation equations, so it is the fixed point of the Bellman evaluation operator. In particular, it is the action-value of the policy $\pi$ on $M_{H_{w},+}$. The argument is identical for $M_{w,-}$ with $Q(s,a)=-\phi(s,a)^{T}w$.
F.4 Proof of Theorem 4.5
Consider the MDP class described in Appendix F.1 and Appendix F.2. First, we know from Lemma F.1 that all instances of the BPI and PE problem characterised by the MDP class $\mathcal{M}$ and target policies $\Pi$ satisfy Assumption 4.2 ($Q^{\pi}$ is realizable for every $\pi$) with the feature map $\phi(\cdot,\cdot)$ defined in Appendix F.1. The proof follows the same reasoning as in Appendix E.4. We consider policy-free queries.
@@
The learner is sample-efficient so $n$ is polynomial in $d$. Specifically, there exists some constant $\alpha>0$ and some integer $T$ such that $n\leq\alpha d^{T}$. Consider $N\in\mathbb{N}$ s.t. $2^{N}\leq d<2^{N+1}$ and set $d_{+}=2^{N}$. Note that $d_{+}\geq d/2$. Fix $W=\exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{K}})$. If $n\leq W$, we can show the learner cannot solve the PE or BPI problems. Consider both cases:
F.4.1 Case 1
Suppose $W<n$, then following the same steps as in E.4.1,
$W<n$ $\implies K\geq c^{\prime}\log\log d\qquad\text{ for a constant $c^{\prime}>0$ and $d$ sufficiently large.}$. (4)
F.4.2 Case 2
$n\leq W\implies n_{k}\leq\exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{k}})$ for $k=1,...,K$. Using the same steps as in E.4.2 in the proof of Theorem 4.4 but with $\pi_{H_{w}}(s^{+}(a))$ replaced by $s^{+}_{H_{w}}(a)$, we can inductively define the sequence of nested subspaces $B_{K}\subset...\subset B_{1}$ characterising $w$ ($\in B_{K}$) and used in the successor state function $s^{+}_{M}$ s.t. for $k=1,...,K$, $\dim B_{k}=2^{\lceil N/4^{k}\rceil}\geq 2^{8}$ and
@@ F.4.2 Case 2
$V^{\star}_{M_{H_{w},-}}(\bar{s})=Q^{\star}_{M_{H_{w},-}}(\bar{s},w)=-\phi(\bar{s},w)^{T}w=-w^{T}w=-1.$
If the learner outputs a policy taking action $a$ s.t. $a^{T}w>0$, it will incur an error greater than $1$ on $M_{H_{w},-}$ and similarly for $M_{H_{w},+}$ if it produces an action $a$ s.t. $a^{T}w\leq 0$. Even if it randomizes between both, with probability at least $1/2$ it will incur an error of at least $1$ on one of the MDPs.
@@
Appendix G Proofs for Fully-Adaptive Setting
G.1 Proof of Theorem B.2
@@ Appendix G Proofs for Fully-Adaptive Setting
Step 1: The learner selects an arbitrary query $(s_{1},a_{1})$ s.t. $\|\phi(s_{1},a_{1})\|_{2}=1$ (possible by Assumption B.1). The learner receives from the environment the reward $r_{M}(s_{1},a_{1})$, the transition function $p_{M}(\cdot|s_{1},a_{1})$ and evaluations of the target policy $\pi_{M}$ for all states in the support of the transition function $p_{M}(\cdot|s_{1},a_{1})$.
@@
The feedback to the learner up to round $k$ means the learner has access to $v_{1},...,v_{k-1}$. This together with Assumption B.1 ensures the query-choice is possible.
@@
$\exists(\alpha_{i})_{i=1}^{k}\quad\text{ s.t.\ one of them is non-zero and }\qquad\sum\limits_{i=1}^{k}\alpha_{i}v_{i}=0.$
Let $j$ be the largest index s.t. $\alpha_{j}\neq 0$, then
@@
G.2 Proof of Theorem B.3
G.2.1 MDP Class
We consider the same MDP class as for Theorem F - see Appendix F.1.
G.2.2 Instance of the class
@@
s.t. $\dim B_{k}=d-k$. Since $\dim B_{d-1}=1$, there is some $w\in B_{K}\cap\partial\mathcal{B}$ s.t. $B_{d-1}=\langle w\rangle$. Let $H_{w}=\{B_{1},...,B_{d-1}\}$ denote the set of nested subspaces. Every MDP $M\in\mathcal{M}$ is fully characterised by the sequence of subspaces $H_{w}$ and a sign $+$ or --. Hence, they are denoted by $M_{H_{w},+}$ or $M_{H_{w},-}$.
@@
See Appendix C for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$.
Transition Function: The transition function for an MDP $M\in\mathcal{M}$ depends on the sequence of nested subspaces $H_{w}$ but not on the sign $\pm$. The transition function is therefore the same for $M_{H_{w},+}$ and $M_{H_{w},-}$. If $A$ is a subspace of $\mathbb{R}^{d}$, let $\text{proj}_{A}(x)$ denote the orthogonal projection of $x$ onto $A$. See Appendix C for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$ and sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. The successor state of a state-action pair $(s,a)$ is deterministic and only depends on the chosen action (and not the current state), so we will denote the unique successor state when taking action $a$ by $s^{+}_{H_{w}}(a)=a$, which is defined as:
@@
This is defined in the same way as in the proof of Theorem F - we refer the reader to Appendix F.1 for an explanation of why this definition is well defined.
@@
G.2.3 Proof of Theorem B.3
Consider the MDP class described in Appendix G.2.1 and Appendix G.2.2. First, we know from Lemma F.1 that all instances of the BPI and PE problem characterised by the MDP class $\mathcal{M}$ and target policies $\Pi$ satisfy Assumption 4.2 ($Q^{\pi}$ is realizable for every $\pi$) with the feature map $\phi(\cdot,\cdot)$ defined in Appendix F.1.
@@
In Appendix G.2.2, we fixed $\dim B_{k}=d-k$. If $a_{1},...,a_{k}$ are not linearly independent, the dimension of $B_{k}$ may be greater than $d-k$. In this case, we just restrict $B_{k}$ arbitrarily such that the sequence remains nested. In particular, we have $\dim B_{d-1}=1$ and $B_{d-1}=\langle w\rangle$ for some $w\in\partial\mathcal{B}$. Let $H_{w}=\{B_{1},...,B_{d-1}\}$. Consider the BPI and PE problem instances with MDPs $M_{H_{w},+}$ and $M_{H_{w},-}$. The transition function is the same on $M_{H_{w},+}$ and $M_{H_{w},-}$.
@@
If the learner outputs a policy taking action $a$ s.t. $a^{T}w>0$, it will incur an error greater than $1$ on $M_{H_{w},-}$ and similarly for $M_{H_{w},+}$ if it produces an action $a$ s.t. $a^{T}w\leq 0$. Even if it randomizes between both, with probability at least $1/2$ it will incur an error of at least $1$ on one of the MDPs.